# Predictive AI with External Knowledge Infusion for Stocks

### v1

@@ 5.2 Hyperparameters
Table 7: Comparison of TA-HKGE without various components of the architecture as described in Section 5.4.
@@
5.3 Baselines
First, we consider a simple (TRANSF) (Vaswani et al., 2017) acting on the historical price information $\mathcal{P}$. Next, we consider the Spatio-Temporal Hypergraph Convolution Network (STHGCN) (Sawhney et al., 2020), which uses a gated temporal convolution over hypergraphs constructed from the stock asset sector classifications, and the Spatio-Temporal Hypergraph Attention Network (STHAN-SR) (Sawhney et al., 2021), which models stock dependencies via hypergraphs to capture higher-order relations, with a hyperedge indicating the presence of a first-order or second-order relation. Finally, we consider GCNKG, a two-layer GCN that considers a homogenous directed graph version of the current time snapshot of the constructed knowledge graphs to obtain node embeddings. These are then concatenated with sequential embeddings and used to calculate the prediction scores. The loss function used for training the baselines is the linear combination of pair-wise ranking loss (Equation 4) and direction loss (Equation 7) defined in Appendix B. All the hyperparameters used for training the baseline and our model are the same.
5.4 Ablations and Additional Experiments
How much does the sequential embedding as asset features and modeling heterogeneous entities help in model performance?
We wish to study how much the components of our method affect the model performance. To that end, we empirically compare multiple variants of our model TA-HKGE without Temporal point process. For the first ablation WOTPP, we remove the embeddings obtained from the temporal process modeling and only consider sequential and relational embeddings. For the second variant, LSTM*, we replace the transformer-based sequential embedding in WOTPP with an LSTM. Next, we consider WOSEQ, in which the sequential embedding update for the stock assets as input for the relational embedding layer is removed, along with the embeddings obtained from the temporal point process model. Finally, the temporal point process embeddings and the heterogeneous transformation of entity features are removed to obtain WOHK. Table 7 provides detailed results, from which it is clear that each of the components of our architecture plays an important role in improving the performance of the algorithm.
How well the model perform when the training, validation, and testing time in each phase is increased?
It is also essential to see how well the model with temporal point process performs when the training, validation, and testing times of each phase are increased since drifts in the stock price over time and changes in the nature of external events can have a negative effect on the model. To measure the extent of this effect, two more training setups are considered. In the first setup ($750$TR), we consider a training period of 750 trading days (3 years) with 75-day validation and a 200-day testing period. In the next setup (1000TR), the training period is 1000 trading days (4 years), with 100-day validation and a 300-day testing period. Table 4 shows that the performance is affected when the training period is increased. However, the results obtained from the model still beat risk-free returns as indicated by the Sharpe ratio.
Table 8: Best and Worst Case ROI Scenarios
$\Delta$ Dataset Top 1 (IRR) Top 5 (IRR)
Best Worst Best Worst
1 N100 4.06 -3.65 3.67 -3.26
S&P 2.98 -2.79 2.65 -2.41
NIFTY 5.12 -3.41 3.77 -3.43
5 N100 4.60 -3.10 4.14 -2.89
S&P 3.62 -2.66 2.78 -2.27
NIFTY 4.69 -2.91 4.52 -3.09
20 N100 4.26 -3.22 4.80 -2.21
S&P 4.68 -1.58 4.33 -1.12
NIFTY 6.13 -2.56 5.61 -2.32
Since, in practice, it is a difficult task to find the exact value of the closing price of a stock, the best and worst-case scenario in terms of IRR is calculated as well. The best case scenario is when the stock is bought at its low price on day $T$ and sold at its high price on day $T+\Delta$, whereas the worst case is when the stock is bought at its high price on day $T$ and sold at its low price on day $T+\Delta$. Table 8 shows the obtained best and worst case IRR using our model, including the Temporal point process. It is observed that the variance in the best and worst IRR decreases as the holding period increases. Also, it is to be noted that accurately predicting the closing price is also an important task for real-life applications of this work.
What is the standard deviation of metrics across the 24 phases?
The market can be bullish or bearish during certain phases, so it is important to look at the standard deviation of the obtained metrics across all the phases. Table 6 shows the standard deviation of metrics across all 24 phases obtained using our model with a Temporal point process. The standard deviation is comparatively lower for the larger datasets S&P500 and NIFTY500. Also, it can be noticed that the standard deviation is lower when the number of stocks selected is higher.
6 Conclusion
In this paper, for the first time, we try to address the problem of forecasting under external influence. To set benchmarks we built Stock Market Knowledge Graphs ($\mathsf{SmKG}$) that offer a systematic framework for constructing knowledge graphs encompassing comprehensive information about publicly available share market data. Our knowledge graphs are designed to capture the temporal and heterogeneous nature of the data and are specifically constructed for two real-world share markets, namely NASDAQ and NIFTY. By integrating various sources of information, including general news events, financial statements and ratios, macroeconomic indicators, corporate events, inter-stock relationships, and sector classification, our framework provides a unified and consolidated view of the share market.
Furthermore, we introduce a model called TA-HKGE, which leverages the power of heterogeneous graph attention networks, temporal process models, and time-aware translation-based knowledge graph embedding techniques to effectively learn the underlying structures and relationships driving stock prices using the constructed knowledge graph. We empirically demonstrate the validity and effectiveness of our proposed methods on the task of stock ranking based on relevant metrics and showcase the importance and necessity of the various components incorporated in our framework. Overall, our research contributes to the field of share market analysis by providing a comprehensive and efficient framework to solve a forecasting problem by incorporating external knowledge in the form of a temporal knowledge graph.
Appendix
Appendix A Loss Functions
A.1 KGE Loss Function
The time-aware knowledge graph embeddings are learned using a translation-based model type loss function, in which the element-wise addition of head entity, relation embedding, and temporal embedding should be close to the tail embedding.
$\mathcal{L}_{1}(h,r,t,\tau)=\|e^{n}_{h}+e^{r}+e^{\tau}-e^{n}_{t}\|$ (3)
A.2 Listwise Ranking
Earlier works (Feng et al., 2019; Sawhney et al., 2020, 2021) employ the following pairwise loss function as the objective.
$\sum_{i=0}^{N}\sum_{j=0}^{N}\max\left\{0,-\left(\hat{Y}_{i}^{\tau+\Delta}-\hat{Y}_{j}^{\tau+\Delta}\right)\left(Y_{i}^{\tau+\Delta}-Y_{j}^{\tau+\Delta}\right)\right\},$ (4)
where $\hat{Y}_{i}^{\tau+\Delta}$ is the predicted ranking score and $Y_{i}^{\tau+\Delta}$ is the actual ranking score for stock $i$. This approach computes the loss for each pair of scores separately, which might be too loose as an approximation of the performance measures of NDCG and MAP. By contrast, the listwise loss function ApproxNDCG (Bruch et al., 2019) used in the listwise approach can better represent the performance measures. The listwise ranking loss function is given by,
$\text{where ~{}rank}_{i}$ $=1+\sum_{j\neq i}\frac{1}{1+\exp\left(\frac{-(y_{j}-y_{i})}{\text{temperature}}\right)}.$ (6)
Here, $y$ is the predicted ranking score and $s$ is the actual score, which is the IRR for each stock from day $\tau$ to day $\tau+\Delta$.
A.3 Cross-Entropy Losses
Let $D^{\tau+\Delta}_{i}$ be the binary movement direction from the closing price of stock $i$ from day $\tau$ to $\tau+\Delta$, in which 1 denotes rise, and 0 denotes fall, and it is expressed as $D^{\tau+\Delta}_{i}=1(p^{i}_{\tau+\Delta}\geq p^{i}_{\tau})$ where, $p^{i}_{\tau}$ is the closing price of stock $i$ on the day $\tau$. The movement ranking loss is defined as the binary cross entropy between $\hat{Y}^{\tau+\Delta}_{i}$ and $D^{\tau+\Delta}_{i}$.
$\mathcal{L}_{3}(\{\hat{Y}\},\{D\})=-[D\log(\hat{Y})+(1-D)\log(1-\hat{Y})]$ (7)
Also, another cross-entropy loss is calculated between the predicted ranking score and stocks that appeared in the top $K$ rank. Such that $R^{\tau+\Delta}_{i}=1(IRR_{i}\text{is in Top 5})$ is the truth labels.
$\mathcal{L}_{4}(\{\hat{Y}\},\{R\})=-[R\log(\hat{Y})+(1-R)\log(1-\hat{Y})]$ (8)
Appendix B Evaluation Metrics
B.1 Investment Return Ratio (IRR)
The Cumulative Investment Return Ratio (IRR) is defined as:
IRR=Final ValueInitial ValueInitial $\text{IRR}=\frac{\text{Final Value}-\text{Initial Value}}{\text{Initial Value}}\times 100,$
the percentage difference between the value of the investment after and at the start of the holding period.
B.2 Annualized Investment Return Ratio (AIRR)
The Annualized Investment Return Ratio (AIRR) is defined as
$\text{AIRR}=(1+\text{ROI})^{\frac{252}{\Delta}}-1,$ (9)
where $252$ is the average number of trading days in a year, $\Delta$ is the number of holding days, and ROI is the Return On Investment defined as IRR$/100$. The AIRR quantifies the average annual rate of return achieved on the investment, taking into account the compounding effect over time.
B.3 Sharpe Ratio (SR)
SR $=\frac{R_{p}-R_{f}}{\sigma_{p}}$ (10)
where, $R_{p}$ represents the average return of the investment, $R_{f}$ represents the risk-free rate of return, and $\sigma_{p}$ represents the standard deviation of the investment’s returns. The Sharpe ratio quantifies the excess return earned per unit of risk taken. It is commonly used to assess the attractiveness of an investment opportunity based on its risk-adjusted performance. The risk-free rates are calculated for each phase by averaging the 6-month treasury-bill yield obtained from “www.treasurydirect.gov” for the USA and “www.investing.com” for India.
B.4 Normalized Discounted Cumulative Gain (NDCG)
The Discounted Cumulative Gain (DCG) is a metric used to evaluate the quality of a ranked list. It can be calculated as
@@ B.4 Normalized Discounted Cumulative Gain (NDCG)
where, $\text{rel}_{i}$ represents the relevance score of the item at position $i$ in the ranked list. The DCG assigns higher weights to more relevant items at the top of the list by dividing the relevance score by the logarithm of the position. The DCG accumulates these weighted scores up to position $k$ to measure the overall quality of the ranked list. The Ideal Discounted Cumulative Gain (IDCG) is a reference value used to evaluate the quality of a ranked list, where $\text{rel}_{i}$ represents the relevance score of the item at position $i$ in the ideal ranking order. The Normalized Discounted Cumulative Gain (NDCG) is a metric used to evaluate the quality of a ranked list. It can be calculated using the following formula:
[…]
The NDCG provides a normalized measure of the ranking performance, ranging from 0 to 1, with higher values indicating better ranking quality.
B.5 Accuracy (ACC)
The accuracy represents the ratio of correct predictions to the total number of predictions made by the model. It provides an overall measure of the model’s ability to correctly identify the top $k$ stocks at each time point.
[…]

### v2

@@ C.1 Hyperparameters
Table 11: Ablation analysis: comparison of TA-HKGE without various components of the architecture as described in Section 7.
@@
Table 12: Performance of the TA-HKGE model with a 750-day training period (750TR) and a 1000-day training period (1000TR).
$\Delta$ Model Dataset Top 1 Top 5
IRR AIRR SR NDCG ACC IRR AIRR SR NDCG ACC
1 750TR N100 0.112 32.47 0.652 0.268 2.5 0.121 35.67 0.428 0.375 12.12
S&P 0.004 0.92 0.179 0.154 0.25 0.066 18.02 0.306 0.225 2.62
NIFTY 0.139 41.83 -499.2 0.163 0.57 0.094 26.7 0.470 0.233 1.97
S&P 0.158 48.78 0.463 0.162 1.16 0.078 21.81 0.329 0.233 2.15
NIFTY 0.087 24.38 0.515 0.155 0.17 0.111 32.4 0.485 0.227 1.83
5 750TR N100 0.38 21.58 1.043 0.275 2.49 0.617 36.34 1.035 0.384 12.46
S&P 0.718 43.42 1.923 0.166 1.05 0.457 25.85 0.856 0.236 3.39
NIFTY 0.447 25.2 1.008 0.159 0.5 0.332 18.21 0.665 0.231 1.93
1000TR N100 0.411 22.97 0.984 0.279 3.27 0.645 38.25 1.006 0.390 13.81
S&P 0.142 7.4 0.597 0.169 1.03 0.621 36.65 1.433 0.244 4.14
NIFTY 0.901 57.1 2.102 0.157 0.71 0.638 37.76 1.181 0.23 2.16
20 750TR N100 0.378 4.87 1.536 0.258 2.17 1.456 19.98 1.698 0.3723 11.58
S&P 0.588 7.66 1.254 0.187 1.35 1.256 17.03 1.420 0.258 4.34
NIFTY 2.637 38.8 3.404 0.162 0.1 2.332 33.7 2.107 0.235 2.51
1000TR N100 1.334 18.18 1.433 0.265 4.01 2.082 29.64 2.440 0.373 11.35
S&P 0.907 12.05 1.767 0.162 0.11 1.853 26.02 1.945 0.237 3.6
NIFTY 0.613 8.0 0.565 0.147 0 1.679 23.34 1.724 0.219 1.33
Appendix D Data Description
The dataset consists of three main components present in the Phase-Stock-KG/data/ directory:
  • [country]_relations.json – Contains relationships between entities in the stock market of the USA or India.
  • [country]_nodes.json – Represents the nodes (companies, sectors) in the knowledge graph.
  • [company_ticker]-[name].csv – Contains historical stock price movements, including features Close, Open, Low, and High prices, and trading volume for each trading day.
D.1 [country]_relations.json & [country]_nodes.json
These files contain the relationships between different entities in the stock market and nodes in the knowledge graph. The information required to create these files was derived from multiple sources, including:
  • Financial Reports: Balance sheets, cash flow statements, and financial statements of individual companies.
  • News Articles: Relevant financial news that may influence stock prices and company relationships.
  • Events: Events in India and the USA that could affect market trends.
D.2 [company_ticker]-[name].csv
The directory data/[nifty500/nasdaq100/sp500] contains CSV files with time-series data of stock price indicators such as Open, High, Low, Close, and Volume for each company.
D.3 Knowledge Graph Construction
The temporal_kg_nifty.pkl and temporal_kg.pkl files, for the Indian and US markets, respectively, are generated by processing relationships and nodes extracted from stock market data. Initially, JSON files such as [country]_relations.json and [country]_nodes.json are loaded to extract the relationships between different entities and the corresponding nodes in the knowledge graph. Each node is mapped to its corresponding attributes.
Once the nodes are mapped, relationships (consisting of a head entity, tail entity, relation type, and timestamps) are iterated and linked to the respective entities, forming a structured knowledge graph. The temporal aspect was introduced by incorporating time-based stock market events from CSV files containing financial statements, news articles, and event data from India and the USA. These time-sensitive events are mapped to the corresponding entities in the graph to ensure a dynamic representation of market fluctuations.
(a) NASDAQ
(b) NSE
Figure 9: Visualization of a subgraph from the constructed $\mathsf{STOCKnowledge}$s for (left) NASDAQ and (right) NSE share markets. Entities and relations are highlighted by type.
Handling timestamps:
  • If an expiry timestamp is available, it is used.
  • If only a timestamp is present, expiry is assumed as one day later.
  • If no timestamp exists, it defaults to 1970-01-01.
The processed relations are stored as
[head,tail,relation_id,timestamp,expiry_timestamp].[\;head,\;tail,\;relation\_id,\;timestamp,\;expiry\_timestamp\;]. (14)
The data were sorted by timestamp for chronological ordering. Finally, the processed temporal knowledge graph was serialized and stored as temporal_kg_nifty.pkl and temporal_kg.pkl using the pickle.
D.4 Batch Dataset Creation from Historical CSV and PKL Files
The batch dataset was constructed using historical stock data from CSV files and a temporal knowledge graph from PKL files.
D.4.1 Stock Data from CSV Files
  • Iterate over stock price CSV files stored in the directory (e.g., data/nasdaq100/).
  • Each file corresponds to a company and contains columns: Date, Open, High, Low, Close, Volume.
  • Ignore files with less than $2800$ rows to ensure sufficient training samples.
D.4.2 Time-Series Windowing
For each company:
  • Apply a sliding window approach with window size $W$ and prediction horizon $\Delta$.
  • Normalize input features by dividing by the first value in the window.
  • Predict returns at different horizons: $T=\{1,5,20\}$.
  • Extract low and high prices for multiple horizons.
D.4.3 Loading the temporal knowledge graph
  • Load the preprocessed knowledge graph from temporal_kg.pkl.
  • Extract relevant temporal facts corresponding to each time window.
  • Constructs a sector-based graph using adjacency relations.
D.4.4 Final Dataset Construction
For each time window:
  • Collect stock price window features and targets.
  • Extract corresponding temporal KG snapshot.
  • Form a dataset tuple:
@@ D.4.4 Final Dataset Construction
The final output is a file data/pickle/[dataset]/full_graph_data.pkl containing a dictionary with keys train and company, which consist of a batch dataset containing stock price sequences and the temporal knowledge graph and a dictionary of company IDs, respectively.
D.5 Node Data Format
The nodes JSON file consists of structured data that represent entities in the knowledge graph. Each node entry contains the following:
  • identity: A unique identifier for the node.
  • labels: A list of labels describing the type of the entity (e.g., ”Person”, ”Company”, etc.).
  • properties: Key-value pairs defining attributes of the node, such as name and id
  • elementId: A system-generated unique identifier for the node.
Example JSON node entry:
"n": {
"identity": 0,
"labels": ["Person"],
"properties": {
"name": "Tommy Millner",
"id": 114689399
"elementId": "0"
In this example, a node labeled Person represents Tommy Millner with an assigned ID of 114689399.
D.6 Relation Data Format
In this example, there is a ”SUBSIDIARY” relationship between the nodes with identities 1007 and 2591, with a relationship ID of ”P355”.
The relations JSON file captures the relationships between nodes and describes how entities are linked. Each relationship entry consists of the following:
  • identity: A unique identifier for the relationship.
  • start: The identity of the starting node in the relationship.
  • end: The identity of the ending node in the relationship.
  • type: A string indicating the relationship type (e.g., ”SUBSIDIARY”, ”OWNS”, ”EMPLOYS”, etc.).
  • properties: Additional metadata, such as a unique id for the relationship.
  • elementId: A system-generated identifier for the relationship.
  • startNodeElementId / endNodeElementId: Unique element IDs of the connected nodes.
Example JSON relation entry:
"r": {
"identity": 0,
"start": 1007,
"end": 2591,
"type": "SUBSIDIARY",
"properties": {
"id": "P355"
"elementId": "0",
"startNodeElementId": "1007",
"endNodeElementId": "2591"
This structured data enables efficient querying and analysis within the knowledge graph, forming the basis for graph-based learning and reasoning. The relationships between entities allow the construction of a knowledge-driven system for applications such as recommendation systems, risk analysis, and trend forecasting, among others.
Appendix E Web Scraping Details
E.1 Entity Recognition and Event Extraction for News Articles
Entity Recognition
Entities such as companies, dates, and financial metrics are recognized using predefined patterns and keywords in the news content. The functions earning_pattern and guidance_pattern use regular expressions and keyword matching to identify the relevant entities and events.
Event Extraction
  • Earnings Events: The earning_pattern function looks for keywords like ’beat’, ’miss’, and ’in-line’ within the news content to determine if a company has met, exceeded, or fallen short of earnings expectations. It extracts the relevant information and categorizes the event as ”BEATS_EXPECTATION”, ”MISSES_EXPECTATION”, or ”INLINE_EXPECTATION”.
  • Guidance Events: The guidance_pattern function searches for keywords indicating changes in company guidance, such as ’raise’, ’cut’, or ’in-line’. It categorizes the event as ”RAISES_GUIDANCE”, ”CUTS_GUIDANCE”, or ”INLINE_GUIDANCE”.
Timestamp Extraction
The get_timestamp function extracts timestamps from the news content. It identifies the location of the timestamp within the content structure and converts it to a datetime object using the Pandas library.
Head and Tail Recognition
  • Head: The head of a triple typically represents the subject of the event, such as the company name. It was extracted from the news content using keyword matching and context analysis.
  • Tail: The tail of a triple represents the object or outcome of the event, such as the financial metric or result. It was identified using similar keyword matching and context analysis techniques.
This process ensures that financial events are accurately extracted and structured for further analysis or integrated into a graph database.
E.2 Profile and Relationship
The notebook wikidata_create.ipynb was designed to create a knowledge graph (KG) from Wikidata for companies listed on NASDAQ and NYSE. It involves several steps, including data loading, processing, mapping, and storing the relationships between the companies and their attributes.
CSV files containing mappings between ticker symbols and Wikidata codes for NASDAQ and NYSE companies were loaded.
E.2.1 Ticker List Creation
Iterates through directories containing company data files to create a list of tickers with sufficient data points (more than 2600 rows).
E.2.2 Wikidata Querying
Using pywikibot, the notebook queries Wikidata to fetch the relationships and attributes of the companies. Relationships such as ”OWNED_BY”, ”SUBSIDIARY”, ”PARENT_ORGANISATION”, ”CEO”, etc., are considered.
Appendix F Time Complexity
We analyze the computational complexity of the TA-HKGE in terms of the key problem dimensions. Let $N$ be the number of stock assets, $W$ the historical window length, $d_{s}$ the sequential embedding dimension, $d_{r}$ the relational embedding dimension, $L$ the number of Transformer encoder layers, and $H$ the number of attention heads. Let $|\mathcal{E}_{T}|$ denote the number of temporal KG edges (events) in a snapshot at time $T$, $|\mathcal{E}_{0}|$ the number of static KG edges, $L$ the number of Transformer encoder layers, and $H$ the number of attention heads. Let $d_{tpp}$ be the temporal point-process embedding dimension used by the HPGE component, and $K$ be the number of negative samples per event. Let $|\mathcal{E}_{T}|$ be the number of temporal KG edges (events) in a snapshot at time $T$, and $|\mathcal{E}_{0}|$ is the number of static KG edges. Let $M$ be the number of relation types (edge types) in a heterogeneous KG.
Sequential encoder (Transformer) For a standard full-attention Transformer encoder over a length-$W$ sequence, the per-layer time complexity is $\mathcal{O}(W^{2}d_{s}+Wd_{s}^{2})$. Thus, computing the sequential embeddings for all $N$ assets costs $\mathcal{O}\big(NL(W^{2}d_{s}+Wd_{s}^{2})\big)$.
Temporal process embeddings (HPGE/Hawkes on graphs) Learning temporal process embeddings scales with the number of observed events and sampled negatives. If we process $|\mathcal{E}_{T}|$ temporal events and sample $K$ negatives per event, the dominant term is $\mathcal{O}(|\mathcal{E}_{T}|\,K\,d_{tpp})$ for embedding updates (with additional overhead from neighborhood aggregation depending on the selected history truncation).
F.1 Relational encoder (heterogeneous attention over KG)
For a message-passing layer with attention over edges, the complexity is linear in the number of edges times the embedding dimension. With $L_{g}$ HEAT layers (we use $L_{g}=2$), the per-snapshot complexity is approximately $\mathcal{O}\big(L_{g}(|\mathcal{E}_{0}|+|\mathcal{E}_{T}|)\,d_{r}\big)$, ignoring constant factors from relation-type-specific projections.
F.2 Prediction and loss
The final ranking layer is $\mathcal{O}(Nd)$ where $d(d_{s}+d_{tpp}+d_{r})$ is the concatenated embedding dimension. The listwise ApproxNDCG loss requires pairwise comparisons within a batch; if computed over all $N$ assets at a time point, it contributes $\mathcal{O}(N^{2})$ per time step.
Combined complexity For one time point $T$ (on one snapshot), an end-to-end training step can be summarized as
$\mathcal{O}$ $\Big(NL(W^{2}d_{s}+Wd_{s}^{2})+|\mathcal{E}_{T}|Kd_{tpp}+L_{g}(|\mathcal{E}_{0}|+|\mathcal{E}_{T}|)d_{r}+\mathcal{O}(Nd)+\mathcal{O}(N^{2})\Big),$